OpenStudy (soccer2080):

rewrite the equations in slope intercept form 4x+5y=15 9x+3y=6

jhonyy9 (jhonyy9):

do you know what mean the slope intercept form ?

OpenStudy (vuriffy):

Do you know what the slope intercept form means?*

OpenStudy (vuriffy):

@jhonyy9 A little fuzzy sentence there, buddy.

OpenStudy (soccer2080):

yea like y=2x+4

jhonyy9 (jhonyy9):

soccer exactly so in this style is correct just rewrite thse two equations in this way - ok. ?

OpenStudy (soccer2080):

oh ok

jhonyy9 (jhonyy9):

hope helped you good luck bye bye

OpenStudy (vuriffy):

Put everything but the y on the other side.

OpenStudy (soccer2080):

ok thank you

OpenStudy (vuriffy):

Do you know how?

jhonyy9 (jhonyy9):

np was my pleasure and anytime bye bye

OpenStudy (soccer2080):

um no

OpenStudy (soccer2080):

ok

OpenStudy (vuriffy):

4x+5y=15 Actually, let's subtract 4x first.

OpenStudy (vuriffy):

5y = 15 - 4x.

jhonyy9 (jhonyy9):

4x+5y=15subtract from both sides 4x and divide both sides by 5

OpenStudy (soccer2080):

ok

OpenStudy (vuriffy):

5y = 15 - 4x. divide by 5. y = 3 - 4x/5.

OpenStudy (soccer2080):

ok

OpenStudy (vuriffy):

9x+3y=6 Do the same with this one: subtract 9x.

OpenStudy (vuriffy):

3y=6 - 9x. Divide by 3.

OpenStudy (vuriffy):

y = 2 - 3x.

OpenStudy (soccer2080):

thank you

OpenStudy (vuriffy):

Of course, do you understand it completely?

OpenStudy (soccer2080):

yes

OpenStudy (vuriffy):

Good job!

OpenStudy (complexdotproduct):

5y=15-4x / 5 -> y = -4x/5 + 3 = -0.8x + 3 3y=6-9x / 3 -> y = -3x + 2